# MultiGran-STGCNFog: Towards Accurate and High-Throughput Inference for Multi-Granular Spatiotemporal Traffic Forecasting

### v1

@@ 5 Experiments
5.1 Experimental Setup
Table 2: Heterogeneous clusters configuration
@@ 5.1 Experimental Setup
We conducted experiments using six heterogeneous fog devices with varying capabilities regarding the CPU, memory, and communication bandwidth. Table 1 provides an overview of these devices, which were configured into ten clusters with distinct memberships shown in Table 2, where network communication was enabled through a symmetric local Ethernet. We evaluated the performance of MultiGran-STGCNFog regarding the forecasting accuracy and inference throughput.
Three benchmark datasets, PEMS04, PEMS07, and PEMS08 were used to evaluate MultiGran-STGCN. Forecasting performance was compared against a set of statistical and GCN approaches introduced in Section 2, including HA, LSTM Cui et al. (2018), GRU Agarap (2018), GCRN Seo et al. (2018), Gated-STGCN Yu et al. (2017), GWNET Wu et al. (2019), OGCRNN Guo et al. (2020), HGCN Guo et al. (2021), and ASTGCN Guo et al. (2019). Baseline parameters were configured as reported in their respective works.
We leveraged MAE, MAPE, and RMSE to measure forecasting performance, and inference efficiency was assessed by the pipeline throughput on PEMS04 (samples/second), with network latency randomly perturbed between 10 and 30 ms to better simulate real-world fog environments.
Table 3: Performance on PEMS04, PEMS07, and PEMS08 datasets. Pink/green/yellow marks the best/second-best/third-best performance.
Metric HA LSTM GRU GCRN Gated-STGCN GWNET OGCRNN HGCN ASTGCN MG-single MG-wms MG-wmt MultiGran-STGCN
PEMS04 MAE 28.38 20.26 20.18 20.78 20.36 18.12 19.76 18.29 20.35 19.00 18.96 18.33 18.11
15 mins MAPE 19.99% 13.87% 13.56% 15.78% 15.77% 12.58% 13.93% 13.02% 14.23% 13.74% 13.56% 12.83% 12.45%
RMSE 41.82 31.71 31.61 31.67 31.14 28.83 30.49 29.84 31.92 29.92 29.74 29.08 28.84
MAE 31.77 22.31 22.23 22.14 21.97 18.86 20.47 19.10 20.93 19.84 19.80 19.04 18.66
30 mins MAPE 22.65% 15.39% 14.94% 16.68% 16.79% 13.13% 14.42% 13.67% 14.57% 14.38% 14.15% 13.21% 12.83%
RMSE 46.49 34.54 34.46 33.65 33.44 29.93 31.55 30.06 33.00 31.11 30.90 30.16 29.82
MAE 38.51 26.41 26.33 24.98 25.17 20.06 21.74 20.62 21.87 21.39 20.89 20.19 19.55
60 mins MAPE 28.20% 18.65% 17.93% 18.68% 19.05% 13.98% 15.41% 14.86% 15.10% 15.47% 14.81% 13.92% 13.45%
RMSE 55.76 40.07 40.08 37.69 37.87 31.62 33.37 32.20 34.82 33.23 32.77 31.89 31.31
PEMS07 MAE 32.82 21.79 21.83 22.71 22.07 19.02 20.50 19.18 22.76 20.87 18.65 18.88 18.56
15 mins MAPE 15.04% 9.26% 9.23% 11.59% 11.57% 8.20% 9.10% 8.29% 9.96% 11.29% 8.11% 8.25% 7.99%
RMSE 47.97 33.84 33.76 34.16 32.84 30.14 31.85 30.25 35.42 31.46 29.72 29.83 29.82
MAE 37.03 24.42 24.50 24.51 24.14 20.32 21.54 20.46 23.89 22.33 19.67 20.14 19.52
30 mins MAPE 17.19% 10.39% 10.36% 12.17% 12.40% 8.67% 9.50% 8.77% 10.30% 12.05% 8.51% 8.73% 8.38%
RMSE 53.99 37.56 37.52 36.83 35.95 32.16 33.48 32.27 37.93 33.71 31.60 31.79 31.55
MAE 45.33 29.51 29.61 28.00 28.08 22.37 23.16 22.31 25.32 24.35 20.90 21.84 20.76
60 mins MAPE 21.56% 12.80% 12.67% 13.45% 14.12% 9.42% 10.20% 9.54% 10.78% 13.29% 9.03% 9.49% 8.95%
RMSE 65.74 44.59 44.42 41.75 41.45 35.12 35.19 35.06 41.12 36.58 34.00 34.47 33.81
PEMS08 MAE 23.11 16.13 16.08 16.85 16.97 14.11 16.04 14.37 16.53 15.14 14.80 14.28 14.03
15 mins MAPE 14.46% 10.02% 10.03% 11.94% 13.58% 9.47% 10.66% 9.20% 10.62% 10.43% 10.37% 9.24% 9.34%
RMSE 34.13 24.95 24.89 25.20 24.92 21.65 24.43 21.92 25.24 22.78 22.59 21.95 21.90
MAE 26.08 17.94 17.86 17.59 17.99 14.84 16.77 15.23 17.01 16.10 15.37 15.05 14.49
30 mins MAPE 16.36% 11.06% 11.11% 12.27% 14.14% 9.84% 11.00% 9.62% 10.89% 11.09% 10.80% 9.61% 9.55%
RMSE 38.31 27.76 27.66 26.49 26.80 22.94 25.68 23.35 26.27 24.34 23.73 23.32 22.90
MAE 32.00 21.46 21.38 19.12 20.14 16.02 18.17 16.69 17.61 17.49 16.20 16.28 15.12
60 mins MAPE 20.28% 13.25% 13.47% 13.07% 15.27% 10.41% 11.79% 10.38% 11.26% 12.13% 11.43% 10.24% 9.91%
RMSE 46.50 32.79 32.73 29.97 30.31 24.88 27.87 25.58 27.56 26.48 25.35 25.39 24.18
5.2 Spatiotemporal Traffic Forecasting
Table 3 presents the performance of MultiGran-STGCN, baselines, and ablation variants: MG-single (single spatial and temporal granularity), MG-wms (without multi-spatial granularity), and MG-wmt (without multi-temporal granularity). MultiGran-STGCN consistently outperforms baselines and ablation models across 15, 30, and 60-minute horizons.
@@ 5.2 Spatiotemporal Traffic Forecasting
5.2.1 Ablation Study of Multi-spatiotemporal Scale Modeling
Figure 2: Ablation study: performance across prediction horizons.
Fig. 2 provides a more intuitive performance comparison among ablation variants and the complete MultiGran-STGCN model denoted as -full, to explore the performance gain of multi-scale modeling and multi-granular feature fusion design. On all three datasets, the -full model consistently outperforms other variants, and its relatively modest increase in errors indicates its effectiveness in capturing long-term dependencies, as the prediction horizon extends from 15 to 60 minutes. The enormous gap between the -single variant and other models points out the limitation of relying solely on single-dimensional scales, drawbacks of previous studies. This observation suggests that while -wms and -wmt partially incorporate multi-spatiotemporal feature extraction, the lack of comprehensive integration restricts their forecasting accuracy. Besides, it is also clear that as the prediction horizon increases, the error trajectories begin to diverge. The growing divergence underscores our model’s superiority in capturing long-term correlations and handling error accumulation.
5.3 Distributed pipeline-parallel model inference
5.3.1 Heterogeneous vs. Homogeneous Scheduling
Figure 3: Heterogeneous vs. homogeneous scheduling algorithm.
The scheduling methods of GPipe Huang et al. (2019) and PipeDream Narayanan et al. (2019) were adopted as the baseline against GA-DPHDS in terms of pipeline throughput. GPipe utilized homogeneous accelerators, while PipeDream partitioned DNN operators based on a single GPU’s capabilities. Devices and pipeline sequences were randomized, and each cluster was tested ten times to obtain the average.
As shown in fig. 3, GA-DPHDS outperformed the baseline across all cluster configurations, achieving throughput from 14.03 samples/s (Cluster 8) to 67.14 samples/s (Cluster 4), while the baseline peaked at 33.72 samples/s and dropped to 12.51 samples/s (Cluster 10). In Cluster 4, GA-DPHDS reached 67.14 samples/s, more than doubling the baseline’s 30.68 samples/s, highlighting its efficiency in resource-intensive setups. In Cluster 2, GA-DPHDS improved throughput to 49.59 samples/s compared to the baseline’s 22.08 samples/s. However, in Cluster 8, having only devices B and C, GA-DPHDS exhibited only minor improvements due to comparable device capabilities, consistent with table 4. These findings underscore GA-DPHDS’s effectiveness in optimizing throughput across varied cluster setups.
Table 4: Single device throughput (samples/s)
@@ 5.3 Distributed pipeline-parallel model inference
5.3.2 Exploring the Impact of Device Network Bandwidth
Figure 4: Cluster throughput with varying bandwidth scaling factors.
We explored the impact of device bandwidth on pipeline throughput by scaling, where bandwidth settings presented in Table 1 took the scaling factor of 1.0 as the standard. Fig. 4 shows that insufficient bandwidth (e.g., scaling factor 0.025) constrained intermediate data transfer efficiency, creating a system bottleneck. We found that in such cases, GA-DPHDS tended to allocate layers to the device with the highest computing power to minimize transmission costs, reducing pipeline parallelism. For instance, in Clusters 1–4, all layers were assigned to device A, resulting in low pipeline throughput. As the scaling factor increased, the average throughput improved (black line), with clusters containing more homogeneous devices showing higher sensitivity to bandwidth, while clusters 8–10 were less affected. Once bandwidth became sufficient (e.g., scaling factor 1.5), it ceased to be a bottleneck, and device computational capacity (cpu) emerged as the new limiting factor. These results demonstrate the adaptability of GA-DPHDS under varying bandwidth constraints.
5.3.3 Optimization Ablation
Figure 5: Ablation study for layer execution order optimization.
We examined the impact of layer execution order $O$ on throughput, optimized via the NSGA-II genetic algorithm. Without this optimization, layer sequences were randomly arranged, adhering only to input-output dependencies. As shown in Fig. 5, optimized layer execution significantly improved throughput. In Cluster 3, GA-DPHDS achieved 58.17 samples/s, a 41.04% increase compared to 41.24 samples/s without NSGA-II, marking the highest observed improvement. Similarly, in Cluster 4, with all the fog devices, throughput was improved by 27.4%, demonstrating that optimized execution order enhances pipeline parallelism and system efficiency. These results highlight the criticality of execution order optimization in maximizing pipeline throughput.
6 Conclusion and Future Work
In this work, we propose MultiGran-STGCNFog, an efficient GNN inference system with a novel traffic forecasting model. MultiGran-STGCNFog extracts spatiotemporal features across various spatial and temporal scales, and the distributed pipeline-parallel architecture of it enables high-performance inference throughput leveraging heterogeneous fog devices. Specifically, the dynamic mechanism and multi-granular feature fusion strengthen its capability to capture long-term traffic dependencies. The proposed scheduling algorithm GA-DPHDS, brings significant throughput improvement. In the future, we might explore the efficiency of our multi-granular spatiotemporal data modeling and feature fusion approach on transformer-based traffic forecasting models, and develop delicate work-stealing strategies to further balance pipeline workload during the runtime phase.

### v2

@@ 5. Experiments
[…]
PEMS04 MAE 28.38 20.26 20.18 20.78 20.36 18.12 19.76 18.29 20.35 19.00 18.96 18.33 18.11 3.28%
15 mins MAPE 19.99% 13.87% 13.56% 15.78% 15.77% 12.58% 13.93% 13.02% 14.23% 13.74% 13.56% 12.83% 12.45% 5.37%
RMSE 41.82 31.71 31.61 31.67 31.14 28.83 30.49 29.84 31.92 29.92 29.74 29.08 28.84 2.96%
MAE 31.77 22.31 22.23 22.14 21.97 18.86 20.47 19.10 20.93 19.84 19.80 19.04 18.66 4.19%
30 mins MAPE 22.65% 15.39% 14.94% 16.68% 16.79% 13.13% 14.42% 13.67% 14.57% 14.38% 14.15% 13.21% 12.83% 6.62%
RMSE 46.49 34.54 34.46 33.65 33.44 29.93 31.55 30.06 33.00 31.11 30.90 30.16 29.82 2.27%
MAE 38.51 26.41 26.33 24.98 25.17 20.06 21.74 20.62 21.87 21.39 20.89 20.19 19.55 6.04%
60 mins MAPE 28.20% 18.65% 17.93% 18.68% 19.05% 13.98% 15.41% 14.86% 15.10% 15.47% 14.81% 13.92% 13.45% 8.17%
RMSE 55.76 40.07 40.08 37.69 37.87 31.62 33.37 32.20 34.82 33.23 32.77 31.89 31.31 3.35%
PEMS07 MAE 32.82 21.79 21.83 22.71 22.07 19.02 20.50 19.18 22.76 20.87 18.65 18.88 18.56 5.14%
15 mins MAPE 15.04% 9.26% 9.23% 11.59% 11.57% 8.20% 9.10% 8.29% 9.96% 11.29% 8.11% 8.25% 7.99% 6.33%
RMSE 47.97 33.84 33.76 34.16 32.84 30.14 31.85 30.25 35.42 31.46 29.72 29.83 29.82 3.01%
MAE 37.03 24.42 24.50 24.51 24.14 20.32 21.54 20.46 23.89 22.33 19.67 20.14 19.52 6.03%
30 mins MAPE 17.19% 10.39% 10.36% 12.17% 12.40% 8.67% 9.50% 8.77% 10.30% 12.05% 8.51% 8.73% 8.38% 6.68%
RMSE 53.99 37.56 37.52 36.83 35.95 32.16 33.48 32.27 37.93 33.71 31.60 31.79 31.55 3.33%
MAE 45.33 29.51 29.61 28.00 28.08 22.37 23.16 22.31 25.32 24.35 20.90 21.84 20.76 8.20%
60 mins MAPE 21.56% 12.80% 12.67% 13.45% 14.12% 9.42% 10.20% 9.54% 10.78% 13.29% 9.03% 9.49% 8.95% 7.92%
RMSE 65.74 44.59 44.42 41.75 41.45 35.12 35.19 35.06 41.12 36.58 34.00 34.47 33.81 3.74%
PEMS08 MAE 23.11 16.13 16.08 16.85 16.97 14.11 16.04 14.37 16.53 15.14 14.80 14.28 14.03 5.46%
15 mins MAPE 14.46% 10.02% 10.03% 11.94% 13.58% 9.47% 10.66% 9.20% 10.62% 10.43% 10.37% 9.24% 9.34% 2.34%
RMSE 34.13 24.95 24.89 25.20 24.92 21.65 24.43 21.92 25.24 22.78 22.59 21.95 21.90 3.38%
MAE 26.08 17.94 17.86 17.59 17.99 14.84 16.77 15.23 17.01 16.10 15.37 15.05 14.49 7.19%
30 mins MAPE 16.36% 11.06% 11.11% 12.27% 14.14% 9.84% 11.00% 9.62% 10.89% 11.09% 10.80% 9.61% 9.55% 5.60%
RMSE 38.31 27.76 27.66 26.49 26.80 22.94 25.68 23.35 26.27 24.34 23.73 23.32 22.90 4.54%
MAE 32.00 21.46 21.38 19.12 20.14 16.02 18.17 16.69 17.61 17.49 16.20 16.28 15.12 9.86%
60 mins MAPE 20.28% 13.25% 13.47% 13.07% 15.27% 10.41% 11.79% 10.38% 11.26% 12.13% 11.43% 10.24% 9.91% 7.24%
RMSE 46.50 32.79 32.73 29.97 30.31 24.88 27.87 25.58 27.56 26.48 25.35 25.39 24.18 7.02%
5.1. Experimental Setup
Table 3. Heterogeneous clusters configuration
@@ 5.1. Experimental Setup
We conducted experiments using six heterogeneous fog devices with varying capabilities regarding the CPU, memory, and communication bandwidth. Table 1 provides an overview of used devices, which were configured into ten clusters with distinct memberships shown in Table 3, where network communication was enabled through a symmetric local Ethernet. Performance was evaluated in terms of forecasting accuracy and inference throughput.
Three benchmark datasets, PEMS04, PEMS07, and PEMS08 were utilized to evaluate MultiGran-STGCN, compared against a set of statistical and GCN approaches introduced in Section 2, including HA, LSTM (Cui et al., 2018), GRU (Agarap, 2018), GCRN (Seo et al., 2018), Gated-STGCN (Yu et al., 2017), GWNET (Wu et al., 2019), OGCRNN (Guo et al., 2020), HGCN (Guo et al., 2021), and ASTGCN (Guo et al., 2019). Baselines were configured as reported in their respective works.
We leveraged MAE, MAPE, and RMSE to measure forecasting performance. Since the throughput gain was analytically guaranteed in Section 4.4, inference efficiency was assessed via pipeline throughput on PEMS04, with network latency randomly perturbed between 10 ~ 30 ms to better simulate real-world fog environments.
5.2. Spatiotemporal Traffic Forecasting
Table 2 presents the performance of MultiGran-STGCN, baselines, and ablation variants: MG-single (single spatial and temporal granularity), MG-wms (without multi-spatial granularity), and MG-wmt (without multi-temporal granularity). MultiGran-STGCN consistently outperforms baselines and its variants: across 15, 30, and 60-minute horizons, with up to 9.86% boost over top-3 baselines.
@@ 5.2. Spatiotemporal Traffic Forecasting
5.2.1. Ablation Study of Multi-spatiotemporal Scale Modeling
Figure 4. Ablation study: forecasts across horizons.
[…]
Ablation
Fig. 4 provides a more intuitive performance comparison among ablation variants and the complete MultiGran-STGCN model denoted as -full, to explore the performance gain of multi-scale modeling and multi-granular feature fusion design. On all three datasets, the -full model consistently outperforms other variants, and its relatively modest increase in errors indicates its effectiveness in capturing long-term dependencies, as the prediction horizon extends from 15 to 60 minutes. The enormous gap between the -single variant and other models points out the limitation of relying solely on single-dimensional scales, drawbacks of previous studies. This observation suggests that while -wms and -wmt partially incorporate multi-spatiotemporal feature extraction, the lack of comprehensive integration restricts their forecasting accuracy. Besides, it is also clear that as the horizon increases, the trajectories begin to diverge. The growing divergence underscores our model’s superiority in capturing long-term correlations and handling error accumulation.
5.3. Distributed pipeline-parallel inference
Figure 5. Heterogeneous vs. homogeneous scheduling.
[…]
heterogeneous_homogeneous
5.3.1. Heterogeneous vs. Homogeneous Scheduling
The scheduling methods of GPipe (Huang et al., 2019) and PipeDream (Narayanan et al., 2019) were adopted as the baseline against GA-DPHDS in terms of pipeline throughput. GPipe utilized homogeneous accelerators, while PipeDream partitioned DNN operators based on a single GPU’s capabilities. Devices and pipeline sequences were randomized, and each cluster was tested ten times to obtain the average. As shown in fig. 5, GA-DPHDS outperformed the baseline across all cluster configurations, achieving throughput from 14.03 samples/s (Cluster 8) to 67.14 samples/s (Cluster 4), while the baseline peaked at 33.72 samples/s and dropped to 12.51 samples/s (Cluster 10). In Cluster 4, GA-DPHDS reached 67.14 samples/s, more than doubling the baseline’s 30.68 samples/s, highlighting its efficiency in resource-intensive setups. In Cluster 2, GA-DPHDS improved throughput to 49.59 samples/s compared to 22.08 samples/s of the baseline. However, in Cluster 8, having only devices B and C, GA-DPHDS exhibited only minor improvements due to comparable device capabilities, consistent with table 4. These findings underscore GA-DPHDS’s effectiveness in optimizing throughput across varied cluster setups.
Table 4. Single device throughput (samples/s)
@@ 5.3. Distributed pipeline-parallel inference
5.3.2. Exploring the Impact of Device Network Bandwidth
Figure 6. Cluster throughput with bandwidth scaling factors.
[…]
bandwidth_scaling
We explored the impact of device bandwidth on pipeline throughput by scaling, where bandwidth settings presented in Table 1 took the scaling factor of 1.0 as the standard. Fig. 6 shows that insufficient bandwidth (e.g., scaling factor 0.025) constrained intermediate data transfer efficiency, creating a system bottleneck. We found that in such cases, GA-DPHDS tended to allocate layers to the device with the highest computing power to minimize transmission costs, reducing pipeline parallelism. For instance, in Clusters 1–4, all layers were assigned to device A, resulting in low pipeline throughput. As the scaling factor increased, the average throughput improved (black line), with clusters containing more homogeneous devices showing higher sensitivity to bandwidth, while clusters 8–10 were less affected. Once bandwidth became sufficient (e.g., scaling factor 1.5), it ceased to be a bottleneck, and device computational capacity (cpu) emerged as the new limiting factor. These results demonstrate the adaptability of GA-DPHDS under varying bandwidth.
5.3.3. Optimization Ablation
Figure 7. Ablation study: layer execution order optimization.
[…]
layer execution order optimization
We also examined the impact of layer execution order $O$ on throughput, optimized via the NSGA-II genetic algorithm. Without this optimization, layer sequences were randomly arranged, adhering only to input-output dependencies. As shown in Fig. 7, optimized layer execution significantly improved throughput. In Cluster 3, GA-DPHDS achieved 58.17 samples/s, a 41.04% increase compared to 41.24 samples/s without NSGA-II, marking the highest observed improvement. Similarly, in Cluster 4, with all the fog devices, throughput was improved by 27.4%, demonstrating that optimized execution order enhances pipeline parallelism and system efficiency. These results highlight the criticality of execution order optimization in maximizing pipeline throughput.
6. Conclusion and Future Work
In this work, we propose MultiGran-STGCNFog, an efficient GNN inference system with a novel traffic forecasting model, which extracts spatiotemporal features across various spatial and temporal scales, and the distributed pipeline-parallel architecture of it enables high-performance inference throughput leveraging heterogeneous fog devices. Specifically, the dynamic mechanism and multi-granular feature fusion strengthen its capability to capture long-term traffic dependencies. The proposed scheduling algorithm GA-DPHDS, brings significant throughput improvement. In the future, we might develop more delicate scheduling features such as work-stealing to further balance pipeline workload during runtime.